x=(1/20)(y-1)^2+2
what is the vertex,focus,wat direction is opens,axis of symmetry and directrix

Answer by stanbon(75887)

x=(1/20)(y-1)^2+2
what is the vertex,focus,what direction is opens,axis of symmetry and directrix
------------
h = 2 ; k = 1 ; p = 5
---------
vertex: (2,1)
axis of symmetry: y = 1
opens to the right
directrix: h-p = 2-5 = -3
focus: (h + 5,k) = (7,1)
===============================
Cheers,
Stan H.